answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

When rolling 2-six sided number cubes what is the probability of rolling a nine?

There are 36 permutations of two standard dice. Of these, four have a sum of nine, specifically 3+6, 4+5, 5+4, and 6+3. The probability, then, of rolling a nine on a pair of standard dice is 4 in 36, or 1 in 9, or about 0.1111.

How is the process of determining percentiles different if you were given a sample mean rather than individual raw score?

It is impossible to determine the percentiles if you are given only the sample mean since percentiles are a measure of the spread of the data; the mean gives no information on that.

How many different telephone numbers of 7 digits can be formed if the first cannot be zero?

There are a number of issues with this question. Phone numbers are limited to more than just the first number not being 0. What about 9-1-1? What about 4-1-1, or 5-1-1, or 3-1-1? What about 1-Area-Number? What about 1-0-Carrier-country-1-city-number? What about the fact that some exchanges do not permit a 0 or a 1 as the second digit of the number? What about the fact that the sum of the first and third digits are constrained to not be certain sums? (The last two rules are to prevent older rule-base auto-long-distance dialing detection logic from misdetecting an area code.) What about constraints on the last four digits.

With that said, and with only the constraints imposed by the question, i.e. ignoring these other known constraints, there are 9,000,000 different phone numbers in one exchange.

What is the experimental probability of rolling an even number?

The experimental probability of anything cannot be answered without doing it, because that is what experimental probability is - the probability that results from conducting an experiment, a posteri. This is different than theoretical probability, which can be computed a priori. For instance, the theoretical probability of rolling an even number is 3 in 6, or 1 in 2, or 0.5, but the experimental probability changes every time you run the experiment.

What is the probability of getting less than 10 when you roll a dice?

A die normally has six sides with the numbers 1 to 6 on them, so any roll will be less than 10 and thus the probability of getting less than 10 with a die is 1.

With two normal dice, the sum of the digits on the dice added together ranges from 2 to 12 and the probability of getting less than 10 is the same as 1 minus the probability of getting 10 or more.

There are 36 ways the two dice can fall and 10 can be achieved in 3 ways (4&6, 5&5, 6&4), 11 can be achieved in 2 ways (5&6, 6&5) and 12 in 1 way (6&6).

Thus the probability of getting 10 or more with 2 dice is (3+2+1)/36 = 6/36 = 1/6

So the probability of getting less than 10 is 1-1/6 = 5/6

What is the probability of rolling snake eyes at least once in four separate rolls of a pair of dice?

The probability of rolling snake eyes on any one throw is 1 in 36. The probability of NOT rolling a snake eyes is 35 in 36. The only outcome here which we do not desire is that in which a non-snake eye roll occurs four times in a row, therefore, we take (35/36)^4 to find the probability of NEVER getting a snake eyes, and then subtract that answer from 1 to get the chances for any outcome WITH at least one pair of snake eyes.

The chances are 10.66% of getting at least one snake eyes in four rolls.

When two dice are rolled find the probability of getting a sum of 5 or 6?

There are 36 permutations of two dice. Of these, 9 have a sum of 5 or 6, so the probability of rolling a sum of 5 or 6 on two dice is 9 in 36, or 1 in 9, or about 0.1111.

What is probability of each suit in a deck of cards?

The probability of each suit in a standard deck of cards is 13 in 52, or 1 in 4, or 0.25.

There is 3 bags with 2 marbles one green and one blue what is the probablitly of choosing one blue marble one green marble and then another blue marble?

It is not explicitly stated in the question, but it is assumed that you draw one marble from each bag. In this case, you have unrelated sequential probability, similar to tossing three coins. The answer is 0.53 or 0.125.

What is the z score for 0.71?

This question cannot be answered. You need the mean and standard deviation in order to compute a Z score for a Raw score. Please restate the question.

How do you find the range of standard deviation?

Standard deviation is not a range. It is a single number that represents how much the data is a set of observations varies about the mean. Mathematically it is the square root of the sum of the squares of the differences between the samples and the mean divided by N, or divided by N-1. The decision of N or N-1 depends on whether or not the mean was also calculated with the calculation of the standard deviation. If it was the same calculation, use N-1, it the mean was known beforehand, use N.

For clarification of parenthesis, and use of N or N-1, see the Related Link below.

How many ways can you arrange the letters in the word Rebecca?

Rebecca = 7 letters

7!

is

7 * 6 * 5 * 4* 3 * 2

= 5040 different ways

-----------------------------

What are the odds in favor of drawing a spade from an ordinary deck of cards?

An ordinary deck of cards has 52 carbs. There are 4 suits in each deck and each suit has 13 cards. You have a 1 in 4 chance of drawing a spade.

What is the probability of picking 2 hearts at random from a deck of playing cards?

Chance of picking first heart card is 13/52, picking second right after is 12/51, total chance is 13/52 multiplied by 12/51. This works if we pick one card, don't replace it and then pick another. Here is the probability of 2 Hearts if we pick two cards at the same time. 13 Hearts choose 2/ 52 cards choose 2= 13C2/52C2=78/1326=1/17

What is the probability that a ball on a roulette wheel will land on any number from 1 to 12?

There are 38 slots on a roulette wheel, two of them being "0" and "00", and the remainders being all numbers from one to thirty-six inclusive.

This means that your chances of getting any number from 1 to 12 are 12/38, or 6/19, which is is just a bit more than 31.5%

Which comes up more when flipping a quarter Heads or tails?

Neither. Its a 50:50 probability. In the long run, heads will match tails.

What is the difference between population and sample distribution?

Sampling distribution is the probability distribution of a given sample statistic.

For example, the sample mean. We could take many samples of size k and look at the mean of each of those. The means would form a distribution and that distribution has a mean, a variance and standard deviation.

Now the population only has one mean, so we can't do this.

Population distribution can refer to how some quality of the population is distributed among the population.

The probability of flipping a coin three times and getting two heads is?

There are 8 permutations of flipping a coin 3 times, or of flipping 3 coins one time. They are, with the permutations of two heads bolded...

TTT

TTH

THT

THH

HTT

HTH

HHT

HHH

... thus, the probability of flipping a coin 3 times and getting 2 heads is 3 in 8, or 0.375.

Trending Questions
What is the probability of getting 3 sixes in 5 tosses of a die? What is the indicated z score if a graph depicts a standard normal variance with the mean 0 and standard deviation of 1? Adjust your vehicle to help prevent a dangerous event from occurring or to minimize the consequences of an unavoidable crash.? What numbers are going to win the lottery? What is the probability of 2 heterozygous individuals producing a homozygous recessive offspring? What is a preceding occurrence cause or event that establishes conditions from which conflict is likely to develop? When you roll a 10 sided die about what percent of the time would you expect a multiple of 3 to come up? Two dice are rolled find the probability that the sum of observed values is equal to 13? A woman and her husband both show the normal phenotype for pigmentation but both had one parent who was an albino what is the probability that their first child will be an albino? What does fifth - 9fifth percentile mean? What are the control limits for mean and range? What is the most likely outcome of these events? What is an 0001 probability level? What is the probability of choosing a marble that is blue? What is the probability of being dealt 2 face cards? How many simple random samples are possible of size 40 from a population of size 500? How to find the variance of 36 percent of voters favor a given ballot for groups of 30 voters find the variance for the nunmber who favor the measure? How many heads can you have? What is the experimental probability of landing on 5 if a number cube is tossed twenty times and lands on 1 two times and lands on 5 four times? Given a standard deck of card. What is the probability of drawing a black face card given that the card you have drawn is black?